# Magnetic soliton molecules in binary condensates

### v1

@@ I Introduction
Stabilized by a balance between dispersion and nonlinearity, solitons are localized collective excitations in elongated systems, relevant for a broad range of systems ranging from water in narrow channels [1, 2] to optical fibers [3], plasmas [4] and conducting polymers [5, 6]. Moreover, ultracold Bose-Einstein condensates (BECs) provide an especially appealing platform for the investigation of solitons, due to the clean and highly controllable experimental apparatuses available. As classic examples, BECs with defocusing nonlinearities can support dark solitons [7, 8, 9, 10], while bright solitons may form in the presence of focussing interactions [11, 12].
Multi-component BECs allow one to investigate the rich physics offered by vector solitons, non-dispersive waves spanning multiple coupled components  [^1] It is worth noting that, although vector solitons may not always be true solitons in the strict mathematical sense, they possess key soliton characteristics, and we will refer to them as solitons for consistency with the literature.. For the case of two distinguishable components—a binary condensate—experiments have observed dark-bright [14] and dark-antidark solitons [15, 16, 17, 18, 19], as well as the closely related magnetic solitons [20, 21] (see related theory [22, 23]), which occur in balanced BEC mixtures close to the miscible-immiscible transition.
@@ II.1 System and methods
with characteristic length $l=\sqrt{\hbar/m\omega}$, where the mass $m$ is assumed to be the same for both components. Ansatz (1) is substituted into the two-component Gross-Pitaevskii equation (GPE) [28]. The radial part is then integrated out, and subtracting the radial zero-point energy yields the quasi-one-dimensional (quasi-1D) GPEs
$i\hbar\partial_{t}\psi_{i}=\left(-\frac{\hbar^{2}}{2m}\partial_{x}^{2}+\sum_{j}\frac{g_{ij}n_{j}}{2\pi l^{2}}\right)\psi_{i}\,$, (2)
with linear densities $n_{j}=|\psi_{j}(x)|^{2}$ and interaction strengths $g_{ij}=4\pi a_{ij}\hbar^{2}/m$, where $a_{ij}$ is the $s$-wave scattering length between components $i$ and $j$.
Equation (2) is solved with a 4th-order Runge-Kutta method, and stationary solutions are found using imaginary time evolution. To create zero-velocity solitons, we enforce the phase of the order parameter to exhibit a $\pi$ step function for the appropriate component at the desired position for every imaginary time step, similar to the approach used in Refs. [29, 30] to create solitons in one-component systems. We validate this method by evolving the converged solutions in real time. For both isolated at-rest solitons and bound-state pairs produced at stationary points of their inter-soliton potential, we observe no change during the evolution, confirming their stationary state character. This method is also useful for creating zero-velocity soliton pairs that are not necessarily stationary states. For example, the initial state ($t=0$) in Fig. 1 (a,b) was created in this way.
@@ II.2 Parameters and considerations for experiments
While our results are given in dimensionless units, they can easily be translated into SI units. For example, using the mass of 87Rb, our computational parameters can correspond to $n_{0}=50\,\mu m^{-1}$, $\omega=2\pi\times 4969\,$Hz, and $(a_{11},\,a_{12},\,a_{22})=(100,95,100)\,a_{0}$, where $a_{0}$ is the Bohr radius. Alternatively, with the choice of 23Na, our computational parameters can correspond to $n_{0}=200\,\mu m^{-1}$, $\omega=2\pi\times 8611\,$Hz, and $(a_{11},\,a_{12},\,a_{22})=(54.54,51.81,54.54)\,a_{0}$. These interaction combinations are close to the regimes already realized in magnetic soliton experiments, e.g., see [15, 20], and we have checked that our results are qualitatively representative of a broad range of density and trapping strength combinations.
@@ III Formulation of the bound state dissociation criterion
Prior work [22] revealed a travelling-wave solution, the so-called magnetic soliton, of the form $\theta(\zeta,\tau)\equiv\theta(\zeta-U\tau)$ and $\phi_{A}(\zeta,\tau)\equiv\phi_{A}(\zeta-U\tau)$ where $U=v/c_{s}$ is the speed of the soliton. The static magnetic soliton $(U=0)$ takes the following simple form
@@
where $\Theta(\zeta)$ is the Heaviside step function. We define the $(+)$ solution as spin-up magnetic soliton, since $s(x)>0$ around the core [see soliton on the right in Fig. 1(a)]. Conversely, the ()(-) solution is referred to as a spin-down soliton. Note that, for $U=0$, the bright component ($n_{i}(x)>n_{0}$ around the core) has no phase variation and the dark component ($n_{i}(x)<n_{0}$) exhibits a $\pi$ phase jump.
An intriguing fact is that a pair of oppositely polarized magnetic solitons, initially cospatial, can be generated with opposite velocities by imprinting a relative-phase domain wall of the form [31]
@@
where $b$ is its characteristic width. The $\tau>0$ dynamics of the pair falls into one of two categories, as can be predicted by the following energy consideration. Using Eq. (4), we can obtain the energy of a single magnetic soliton
$E_{\text{MS}}(U)=n\hbar c_{s}\sqrt{1-U^{2}},$ (7)
@@
$E_{\text{RPDW}}(b)=2n\hbar c_{s}/b.$ (8)
@@
On the other hand, when the initial width is narrow, $b<1$, the energy of the generated soliton pair is too large to form well-separated solitons. Instead, they quickly evolve into a pair of static magnetic solitons ($U=0$) at finite separation, before again moving towards one another, thus forming an oscillatory bound state (analogous to an excited molecule). This counterintuitive behavior may be understood by noting that the energy of a soliton decreases with increasing velocity [see Eq. (7)], which corresponds to a negative effective mass
$m_{\text{eff}}=\frac{1}{Uc_{s}^{2}}\frac{dE_{\text{MS}}}{dU}=-\frac{n\hbar}{c_{s}}\frac{1}{\sqrt{1-U^{2}}}.$ (10)
@@ IV.1 Soliton-soliton interaction potential
Figure 2: (a) Interaction potential between two magnetic solitons with opposite polarization as a function of their separation, calculated using the GPE (blue points) and the uniform-$n$ approximation $V_{\text{uni}}$ (black curve). Inset: Oscillation period of the bound states as a function of their maximum separation. (b) Interaction potential for pairs with like polarization. When $\Delta x\to 0$, the potential approaches the energy of a single magnetic soliton (horizontal dashed line). Magenta circles highlight the two cases shown in Fig. 1(c,d).
@@
$V=\text{sign}(m_{\text{eff}})\left[E_{\text{2MS}}^{\rm GPE}(\Delta x)-E_{0}\right],$ (11)
where $E_{\text{2MS}}^{\rm GPE}$ is the energy of the system for two solitons at rest separated by $\Delta x=x_{2}-x_{1}$, with $x_{i}$ being the position of soliton $i$. The reference energy $E_{0}$ corresponds to the system energy for maximally separated solitons at rest.
@@
Figure 2(a) shows $V$ for two solitons with opposite polarization. The blue points are calculated using the GPE (2) with the approach outlined in Sec. II.1. Importantly, there is a minimum at $\Delta x=0$, indicating a restoring force towards zero separation  [^3] It is worth noting that, since magnetic solitons have a negative effective mass, i.e., $\text{sign}(m_{\text{eff}})=-1$ [see Eq. (10)], they are in fact accelerated towards regions of potential maxima, although these appear as minima in terms of our definition of $V$.. We have confirmed that zero-separation bound states remain stationary under real time evolution of the GPE (not shown). In contrast, Fig. 2(b) shows $V$ for a pair of solitons with like polarization. In the limit $\Delta x\to 0$, $V$ approaches the energy of a single magnetic soliton in the system (horizontal dashed line; see Appendix A for details). To understand this, recall that the wavefunction phase has two $\pi$ jumps in the same component. Taking the limit $\Delta x\to 0$ but excluding $\Delta x=0$, the region where the density is forced to zero approaches a single point, just as for a single soliton. For increasing $|\Delta x|$, $V$ decreases monotonically, indicating that like-sign solitons repel each other and thus cannot form bound states [see Fig. 1(d)].
@@
A semianalytic prediction for the intersoliton potential can be made based on the uniform-$n$ approximation outlined in section III. The potential energy can be obtained using the difference between the energy of a relative-phase domain wall of width $b<1$ and that of two isolated magnetic solitons at zero velocity, i.e.,
$E_{\text{uni}}=E_{\text{RPDW}}(b)-2E_{\text{MS}}(U=0)=2n\hbar c_{s}\left(\frac{1}{b}-1\right).$ (12)
@@
To determine the relation between $b$ and $\Delta x$, we evolve relative phase domain walls under the uniform-$n$ equations (3) and extract $\Delta x$ as the maximum soliton separation at rest after a quarter of an oscillation period. The resulting semianalytic prediction of $V_{\text{uni}}=\text{sign}(m_{\text{eff}})E_{\text{uni}}$ as a function of $\Delta x$ is plotted in Fig. 2(a) as solid lines, demonstrating good agreement with the GPE results for moderate soliton separations. Note, however, that $V_{\text{uni}}$ diverges for $\Delta x\ll\xi_{s}$, because the uniform-$n$ constraint significantly overestimates the kinetic energy when $b\ll 1$.
@@ IV.2 Critical binding energy
Relative-phase domain walls of width $b_{0}$ [Eq. (6)] are first imprinted on otherwise uniform BECs. These are then subjected to a short imaginary-time evolution ($\Delta t=312i/\omega$) to relax the density near the soliton cores, which need not be uniform. The resulting states, which serve as our initial conditions for real-time evolution, differ slightly in width from $b_{0}$; we extract the actual $b$ by fitting their phase profiles [Eq. (6)]. The fit errors are small, with error bars in Fig. 3 generally smaller than the symbols (see Appendix B).
The system energy $E_{\rm RPDW}^{\rm GPE}$ is calculated for each state, with the reference energy of two maximally separated solitons subtracted to obtain
$E=E_{\rm RPDW}^{\rm GPE}(b)-E_{0}.$ (13)
@@
Figure 3 shows the energy of relative phase domain walls [Eq. (13)] as a function of inverse width. Squares (circles) denote soliton pairs that evolve to be unbound (bound) in real-time simulations. Two example simulations are plotted as insets, one on each side of the binding threshold, i.e., $b^{-1}=0.9$ and $b^{-1}=1.1$. The analytic uniform-$n$ prediction [Eq. (12)] agrees quantitatively with the GPE simulations near the threshold. At larger $1/b$ (slow-moving solitons), deviations arise from the constraint of uniform total density (see Sec. IV.1).
@@ IV.3 Collisions involving bound states
Since solitons of like polarization repel and those of opposite polarization attract, an excited bound state behaves like a dipole oscillating with period $T$. The time-dependent polarity of these dipoles is illustrated in Fig. 4 by black arrows, which point from spin-down (red) to spin-up (blue) solitons within each bound state. Figure 4(a) shows a collision between two tightly bound states, each with an initial T=31.04T=31.04\,ms. In this case, the polarities of the two bound states remain opposite throughout the collision, leading to inter-bound-state repulsion and a clear bounce, with both bound states remaining intact.
Figure 4(b) shows two tightly bound states (initially T=34.2T=34.2\,ms) with the same polarity (black arrows pointing in the same direction). In this case, the bound-state dipoles tend to attract each other, yet both bound states nevertheless survive the collision. The attraction between same-polarity bound states has striking consequences for two loosely bound states (initially T=128.9T=128.9\,ms), as shown in Fig. 4(c). Here, the inner solitons—having opposite polarization—not only cross each other but are actually torn from their respective bound states and form a new bound state, while the two outer solitons become unbound. Finally, in the case of a single soliton colliding with a bound state (initially T=36.59T=36.59\,ms), Fig. 4(d) illustrates a type of substitution reaction, in which the newcomer replaces one of the bound state’s solitons and ejects it in the process.

### v2

@@ I Introduction
Stabilized by a balance between dispersion and nonlinearity, solitons are localized collective excitations in elongated systems, relevant for a broad range of systems ranging from water in narrow channels [1, 2] to optical fibers [3], plasmas [4] and conducting polymers [5, 6]. Moreover, ultracold Bose-Einstein condensates (BECs) provide an especially appealing platform for the investigation of solitons, due to the clean and highly controllable experimental apparatuses available. As classic examples, BECs with defocusing nonlinearities can support dark solitons [7, 8, 9, 10], while bright solitons may form in the presence of focusing interactions [11, 12].
Multi-component BECs allow one to investigate the rich physics offered by vector solitons, non-dispersive waves spanning multiple coupled components  [^1] It is worth noting that, although vector solitons may not always be true solitons in the strict mathematical sense, they possess key soliton characteristics, and we will refer to them as solitons for consistency with the literature.. For the case of two distinguishable components [14, 15, 16, 17, 18, 17]—a binary condensate—experiments have observed dark-bright [19] and dark-antidark solitons [20, 21, 22, 23, 24], as well as the closely related magnetic solitons [25, 26] (see related theory [27, 28]), which occur in balanced BEC mixtures close to the miscible-immiscible transition.
@@ II.1 System and methods
with characteristic length $l=\sqrt{\hbar/m\omega}$, where the mass $m$ is assumed to be the same for both components. Ansatz (1) is substituted into the two-component Gross-Pitaevskii equation (GPE) [33]. The radial part is then integrated out, and subtracting the radial zero-point energy yields the quasi-one-dimensional (quasi-1D) GPEs:
$i\hbar\partial_{t}\psi_{1}$ $=$ $\left(-\frac{\hbar^{2}}{2m}\partial_{x}^{2}+\frac{g_{11}n_{1}}{2\pi l^{2}}+\frac{g_{12}n_{2}}{2\pi l^{2}}\right)\psi_{1}\,$,
$i\hbar\partial_{t}\psi_{2}$ $=$ $\left(-\frac{\hbar^{2}}{2m}\partial_{x}^{2}+\frac{g_{12}n_{1}}{2\pi l^{2}}+\frac{g_{22}n_{2}}{2\pi l^{2}}\right)\psi_{2}\,$, (2)
with linear densities $n_{j}=|\psi_{j}(x)|^{2}$ and interaction strengths $g_{ij}=4\pi a_{ij}\hbar^{2}/m$, where $a_{ij}$ is the $s$-wave scattering length between components $i$ and $j$, and $g_{12}=g_{21}$.
Equation (2) is solved with a 4th-order Runge-Kutta method, and stationary solutions are found using imaginary time evolution. To create zero-velocity solitons, we enforce the phase of the order parameter to exhibit a $\pi$ step function for the appropriate component at the desired position for every imaginary time step, similar to the approach used in Refs. [34, 35] to create solitons in one-component systems. We validate this method by evolving the converged solutions in real time. For both isolated at-rest solitons and bound-state pairs produced at stationary points of their inter-soliton potential, we observe no change during the evolution, confirming their stationary state character. This method is also useful for creating zero-velocity soliton pairs that are not necessarily stationary states. For example, the initial state ($t=0$) in Fig. 1(a) and 1(a) was created in this way.
@@ II.2 Parameters and considerations for experiments
While our results are given in dimensionless units, they can easily be translated into SI units. For example, using the mass of 87Rb, our computational parameters can correspond to $n_{0}=50\,\upmu\rm{m}^{-1}$, $\omega=2\pi\times 4969\,$Hz, and $(a_{11},\,a_{12},\,a_{22})=(100,95,100)\,a_{0}$, where $a_{0}$ is the Bohr radius. Alternatively, with the choice of 23Na, our computational parameters can correspond to $n_{0}=200\,\upmu\rm{m}^{-1}$, $\omega=2\pi\times 8611\,$Hz, and $(a_{11},\,a_{12},\,a_{22})=(54.54,51.81,54.54)\,a_{0}$. These interaction combinations are close to the regimes already realized in magnetic soliton experiments, e.g., see [20, 25], and we have checked that our results are qualitatively representative of a broad range of density and trapping strength combinations.
@@ III Formulation of the bound state dissociation criterion
Prior work [27] revealed a traveling-wave solution, the so-called magnetic soliton, of the form $\theta(\zeta,\tau)\equiv\theta(\zeta-U\tau)$ and $\phi_{A}(\zeta,\tau)\equiv\phi_{A}(\zeta-U\tau)$ where $U=v/c_{s}$ is the speed of the soliton. The static magnetic soliton $(U=0)$ takes the following simple form
@@
where $\Theta(\zeta)$ is the Heaviside step function. We define the $(+)$ solution as spin-up magnetic soliton, since $s(x)>0$ around the core [see soliton on the right in Fig. 1(a)]. Conversely, the ()(-) solution is referred to as a spin-down soliton. Note that, for $U=0$, the antidark component ($n_{i}(x)>n_{0}$ around the core) has no phase variation and the dark component ($n_{i}(x)<n_{0}$) exhibits a $\pi$ phase jump.
Figure 2: Relative-phase domain walls showing the phase profiles of component 1 (2) as solid (dot-dashed) curves for three different widths, $b=\{0.5,1,2\}$, from dark to light.
An intriguing fact is that a pair of oppositely polarized magnetic solitons, initially cospatial, can be generated with opposite velocities by imprinting a relative-phase domain wall of the form [36]:
@@
where $b$ is its characteristic width. Although the spin density is spatially uniform, $\phi_{A}$ can be symmetrically distributed between the two components, revealing the soliton pairs shown in Fig. 2, where three different widths are considered. The $\tau>0$ dynamics of the pair falls into one of two categories, as can be predicted by the following energy consideration. Using Eq. (4), we can obtain the energy of a single magnetic soliton
$E_{\text{MS}}^{\text{uni}}(U)=n\hbar c_{s}\sqrt{1-U^{2}},$ (7)
@@
$E_{\text{RPDW}}^{\text{uni}}(b)=2n\hbar c_{s}/b.$ (8)
@@
On the other hand, when the initial width is narrow, $b<1$, the energy of the generated soliton pair is too large to form well-separated solitons. Instead, they quickly evolve into a pair of static magnetic solitons ($U=0$) at finite separation, before again moving towards one another, thus forming an oscillatory bound state (analogous to an excited molecule). This counterintuitive behavior may be understood by noting that the energy of a soliton decreases with increasing velocity [see Eq. (7)], which corresponds to a negative effective mass  [^3] For comparison, it is interesting to note that while dark solitons also have negative effective masses [39], certain dark-bright solitons have been predicted to dynamically oscillate between negative and positive effective mass [40].
$m_{\text{eff}}=\frac{1}{Uc_{s}^{2}}\frac{dE_{\text{MS}}^{\text{uni}}}{dU}=-\frac{n\hbar}{c_{s}}\frac{1}{\sqrt{1-U^{2}}}.$ (10)
@@ IV.1 Soliton-soliton interaction potential
Figure 3: (a) Interaction potential between two magnetic solitons with opposite polarization as a function of their separation, calculated using the GPEs (blue points) and the uniform-$n$ approximation $V^{\text{uni}}$ (black curve). Inset shows the oscillation period of the bound states as a function of their maximum separation. (b) Interaction potential for pairs with like polarization. When $\Delta x\to 0$, the potential approaches the energy of a single magnetic soliton (horizontal dashed line). Magenta circles highlight the two cases shown in Fig. 1(c,d).
@@
$V=\text{sign}(m_{\text{eff}})\left[E_{\text{2MS}}(\Delta x)-E_{0}\right],$ (11)
where $E_{\text{2MS}}$ is the total system energy for two solitons at rest separated by $\Delta x=x_{2}-x_{1}$, with $x_{i}$ denoting the position of soliton $i$. The reference energy $E_{0}$ corresponds to the total system energy for maximally separated solitons at rest, so that $V\to 0$ when the solitons are far apart. The sign prefactor in Eq. (11) is included to make the relationship between the force and acceleration more intuitive. Since magnetic solitons have a negative effective mass, $\text{sign}(m_{\text{eff}})=-1$ [see Eq. (10)], the acceleration $a=F/m_{\text{eff}}$ is actually opposite in direction to the force $F$. However, this sign difference is absorbed into our definition of $V$ and bound states thus occur within potential minima.
@@
Figure 3(a) shows $V$ for two solitons with opposite polarizations. The blue points are calculated using the GPEs following the approach outlined in Sec. II.1. Importantly, there is a minimum at $\Delta x=0$, indicating that the solitons experience an attraction toward zero separation. This behavior is qualitatively different from that of real molecular potentials, where the minima occur at finite separation. We have confirmed that zero-separation bound states remain stationary under real time evolution of the GPEs (not shown). In contrast, Fig. 3(b) shows $V$ for a pair of solitons with like polarization. In the limit $\Delta x\to 0$, $V$ approaches the energy of a single magnetic soliton in the system (horizontal dashed line; see Appendix A for details). To understand this, recall that the wavefunction phase has two $\pi$ jumps in the same component. Taking the limit $\Delta x\to 0$ but excluding $\Delta x=0$, the region where the density is forced to zero approaches a single point, just as for a single soliton. For increasing $|\Delta x|$, $V$ decreases monotonically, indicating that like-sign solitons repel each other and thus cannot form bound states [see Fig. 1(d)].
@@
A semianalytic prediction for the intersoliton potential can be made based on the uniform-$n$ approximation outlined in Sec. III. The soliton interaction energy is obtained from the difference between the energy of a relative-phase domain wall of width $b<1$ and that of two isolated magnetic solitons at zero velocity, i.e.,
$E^{\text{uni}}=E_{\text{RPDW}}^{\text{uni}}(b)-2E_{\text{MS}}^{\text{uni}}(U=0)=2n\hbar c_{s}\left(\frac{1}{b}-1\right).$ (12)
@@
The energy of the relative phase domain wall is equivalent to that of two solitons at rest separated by $\Delta x$, provided that $b<1$, because these two configurations correspond to different phases of the same oscillation. To determine the relation between $b$ and $\Delta x$, we evolve relative phase domain walls under the uniform-$n$ equations (3) and extract $\Delta x$ as the maximum soliton separation at rest after a quarter of an oscillation period. The resulting semianalytic prediction of the potential $V^{\text{uni}}=\text{sign}(m_{\text{eff}})E^{\text{uni}}$ as a function of $\Delta x$ is plotted in Fig. 3(a) as solid lines, demonstrating good agreement with the GPE results for moderate soliton separations. Note, however, that $V^{\text{uni}}$ diverges for $\Delta x\ll\xi_{s}$, because the uniform-$n$ constraint significantly overestimates the kinetic energy when $b\ll 1$.
@@ IV.2 Critical binding energy
Relative-phase domain walls of width $b_{0}$ [Eq. (6)] are first imprinted on otherwise uniform BECs. These are then subjected to a short imaginary-time evolution ($\Delta t=312i/\omega$) to relax the density near the soliton cores, which need not be uniform. The resulting states, which serve as our initial conditions for real-time evolution, differ slightly in width from $b_{0}$; we extract the actual $b$ by fitting their phase profiles [Eq. (6)]. The fit errors are small, with error bars in Fig. 4 generally smaller than the symbols (see Appendix B). The system energy $E_{\rm RPDW}$ is calculated for each state, with the reference energy of two maximally separated solitons subtracted to obtain
$E=E_{\rm RPDW}(b)-E_{0}.$ (13)
@@
Figure 4 shows the energy of relative phase domain walls [Eq. (13)] as a function of inverse width. Squares (circles) denote soliton pairs that evolve to be unbound (bound) in real-time simulations. Two example simulations are plotted as insets, one on each side of the binding threshold, i.e., $b^{-1}=0.9$ and $b^{-1}=1.1$. The analytic uniform-$n$ prediction [Eq. (12)] agrees quantitatively with the GPE simulations [Eq. (13)] near the threshold. At larger $1/b$ (slow-moving solitons), deviations arise from the constraint of uniform total density (cf. Sec. IV.1).
@@ IV.3 Collisions involving bound states
Since solitons of like polarization repel and those of opposite polarization attract, an excited bound state behaves like a dipole oscillating with period $T$. The time-dependent polarity of these dipoles is illustrated in Fig. 5 by black arrows pointing from spin-down (red) to spin-up (blue) solitons within each bound state. Figure 5(a) shows a collision between two tightly bound states, each with an initial T=31.04T=31.04\,ms. In this case, the polarities of the two bound states remain opposite throughout the collision, leading to inter bound-state repulsion and a clear bounce, with both bound states remaining intact. Evidence for bound state energy loss is seen in the slightly increasing oscillation periods over time. The energy lost is converted to phonons, which are not visible with this color scheme.
Figure 5(b) shows two tightly bound states (initially T=34.2T=34.2\,ms) with the same polarity (black arrows pointing in the same direction). In this case, the bound-state dipoles tend to attract each other, yet both bound states nevertheless survive the collision. This attractive interaction is only weakly visible due to the tight binding and the relatively fast center-of-mass motions of the bound states. By contrast, the attractive interaction between same-polarity bound states has striking consequences for two loosely bound states (initially T=128.9T=128.9\,ms), as shown in Fig. 5(c). Here, the inner solitons—having opposite polarization—not only cross each other but are actually torn from their respective bound states and form a new bound state, while the two outer solitons become unbound. Finally, in the case of a single soliton colliding with a bound state (initially T=36.59T=36.59\,ms), Fig. 5(d) illustrates a type of substitution reaction, in which the newcomer replaces one of the bound state’s solitons and ejects it in the process.